OpenStudy (anonymous):

Cos = √17/17 Find csc. A.√17/4 B. 4 C.√17/17 D.4/17 (√17) E.1/4 F. None of these

OpenStudy (anonymous):



OpenStudy (ybarrap):

use @satellite73 's clue. b = 17 sin \( \theta \) and csc \(\theta\) = 1/sin \(\theta\) That is, use Pythagorean's Theorem to find b then just find the reciprocal .